OpenStudy (yumeybitaa):

Anthony was tracking the increasing number of animals in the zoo. At 9 a.m., there were 56 animals. At 11 a.m., there were 60 animals. If Anthony made the function f(x) = 2x − 38, what would the 2 represent? The number of animals at midnight The rate at which the number of animals was increasing The length of time he recorded for The total change in the number of animals

OpenStudy (anonymous):

2 is the gradient of the function at any point. So for every 1 hour, there are 2 more animals in a zoo (i.e. the number of animals increase at a constant rate of 2 animals / hour)

OpenStudy (anonymous):

its B hun

OpenStudy (yumeybitaa):

thank you so much ..

OpenStudy (anonymous):

np lemme know if u have anymore

OpenStudy (yumeybitaa):

can you help me with one more ?

OpenStudy (anonymous):

yea

OpenStudy (yumeybitaa):

Given the point (3, 4) and the slope of 6, find y when x = 27. 122 14 141 148

OpenStudy (anonymous):

We can use the Point-Slope form of a line: y−yo=m(x−xo) We were given the slope: m=6 And a starting point, (xo,yo)=(3,4) And an x value, x=27

OpenStudy (anonymous):

Plugging in our values gives us, y−yo=m(x−xo) => y−4=6(27−3)

OpenStudy (anonymous):

27 minus 3, y−4=6(24) 6 times 24, y−4=144 Add 4 to each side, y=144+4 148...

OpenStudy (yumeybitaa):

wow that's confusing ...thank you

OpenStudy (anonymous):

aha np u should give m a medal ;)